five times a number decreased by 6 is equal to the number squared? what is the number?
Answer by mananth(16946)

five times a number decreased by 6 is equal to the number squared? what is the number?
...
let number be x
5x-6 = x^2
+6
5x=x^2+6
-5x
x^2-5x+6=0
(x-6)(x+1)=0
x= 6 OR x=-1
...
m.ananth@hotmail.ca